OpenStudy (anonymous):

What is the slope of the line that passes through the points (1, -2) and (3, -5)? A) B) C) D)

OpenStudy (anonymous):

-3/2

OpenStudy (anonymous):

slope = gradient = m = (y-y)/(x-x) m= (-2-(-5)/(1-3) m= -3/2